OpenStudy (anonymous):

simplifiy the following expression #1. 3^2 * 9^-4/3 #2. 4^3*2^-2/3*8^5/2

OpenStudy (anonymous):

#1 1/2187

OpenStudy (anonymous):

\[\frac{3^{2}\times 9^{-4}}{3}=\frac{3^{2}}{3 \times 9^{4}}\]\[\frac{9}{3\times 6561} = \frac{9}{19683}\]\[\frac{1}{2187}\]

OpenStudy (anonymous):

If I deserve medal then give me one

OpenStudy (anonymous):

thanks i thought i already did sorry

OpenStudy (anonymous):

you deserve a kiss you *muah*

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

\[\frac{4^{3}\times2^{-2}}{3\times 8^{5/2}}\]

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

question for the first one. why divide by 3? is it because of the fraction?

OpenStudy (anonymous):

4/3 fraction is mean divi

OpenStudy (anonymous):

second : 4^3*2^-2/3*8^5/2 I 'n sure question ask

OpenStudy (anonymous):

\[(\frac{4^{3}\times 2 ^{-2}}{3 \times 8^{5}})\div 2\] this one more sen

OpenStudy (anonymous):

my last name is Lam too

OpenStudy (anonymous):

u want me try do this, u can repost see some one answer, yes we the same last name

OpenStudy (anonymous):

u r chiness?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

where u born?

OpenStudy (anonymous):

yes it just say simplify for those problems. the first one anser was 3^-2/3 i just dont know how to do it. =(

OpenStudy (anonymous):

i was born in vietnam.. but i speak chinese. what about you?

OpenStudy (anonymous):

\[\frac{3^{-2}}{3}\] -2 negative than bring 3^-2 down made positve and the top will 1 \[\frac{1}{3*3^{2}}\] the same me,I'm chini but horn in VN now I in USA, I hear ,t talk not well , ny parent very good write and speak ,mandarin and chow cho

OpenStudy (anonymous):

can u do the rest?

OpenStudy (anonymous):

my sister the same your name only i not y here name is Lam mi

OpenStudy (anonymous):

I have to go do some erran

OpenStudy (anonymous):

heheh ok thanks! hope to get help from you soon!

OpenStudy (anonymous):

OK bye

OpenStudy (anonymous):

bye